# VideoSAVi: Self-Aligned Video Language Models without Human Supervision

### v2

@@ Abstract
Recent advances in video-large language models (Video-LLMs) have led to significant progress in video understanding. Current preference optimization methods often rely on proprietary APIs or ground-truth captions to generate preference data (i.e., pairs of model outputs ranked based on their quality or alignment with human judgment), which is then used to train models for video-language alignment. This approach is both costly and labor-intensive. To address this limitation, we introduce VideoSAVi (Self-Aligned Video Language Model), a self-training pipeline that enables Video-LLMs to reason over video content without external supervision. Our approach includes a self-critiquing mechanism that identifies reasoning errors in the model’s initial responses and generates improved alternatives, creating preference pairs directly from video content. VideoSAVi then applies Direct Preference Optimization (DPO), which uses the preference data to iteratively train the model, enhancing temporal and spatial reasoning in video understanding. Experiments show that VideoSAVi achieves state-of-the-art performance on MVBench (74.0%) and delivers significant improvements across other benchmarks, including a 3.9% gain on PerceptionTest and a substantial 6.8% improvement on the challenging EgoSchema dataset compared to baseline models. Our model-agnostic approach is computationally efficient, requiring only 32 frames, offering a promising direction for self-aligned video understanding without reliance on external models or annotations.
@@ 1 Introduction
Vision-language models (VLMs) (Liu et al., 2023; Li et al., 2023; Radford et al., 2021; Wang et al., 2024b; Chen et al., 2024; Li et al., 2024a) have made significant strides by integrating visual perception with the reasoning capabilities of large language models (LLMs) (OpenAI, 2024; Dubey et al., 2024; Ouyang et al., 2022). These models excel in interpreting and generating contextually relevant responses through the combination of image encoders and language generation techniques. Building on this foundation, recent video-large language models (Video-LLMs) (Zhang et al., 2023; Lin et al., 2024; Li et al., 2024b; Zhang et al., 2024d; Wang et al., 2022; 2024e) incorporate temporal dimensions, enabling comprehensive video understanding by transforming video frames into tokens that LLMs can process. While Video-LLMs demonstrate impressive capabilities, they typically require vast, high-quality annotated datasets, making them resource-intensive and limiting their scalability.
@@
Instruction tuning has been pivotal in advancing both VLMs and Video-LLMs (Liu et al., 2023; Brown et al., 2020; Xu et al., 2023; Wei et al., 2021; Wang et al., 2024b; Chen et al., 2024; Zhang et al., 2024c), but creating extensive video instruction datasets incurs substantial costs (Deng et al., 2024). Recent efforts to generate large instruction datasets (up to 1.3M pairs) by distilling knowledge from proprietary models like GPT-4V (Zhang et al., 2024d) have yielded only marginal improvements. This dependence on extensive annotated data and proprietary models restricts the adaptability of Video-LLMs, posing a barrier to broader applications.
@@
Alignment-based approaches have shown promise for improving video understanding. LLaVA-Hound (Zhang et al., 2024b) uses Direct Preference Optimization (DPO) (Rafailov et al., 2023) with text-based ranking of preference data generated by proprietary models, while TPO (Li et al., 2025) targets temporal preference optimization by sampling negative examples from video segments not present in the target clip. However, these methods depend heavily on expensive proprietary APIs or ground truth captions, limiting their accessibility and scalability.
@@
This raises a critical research question: How can we train Video-LLMs to generate high-quality outputs without relying on proprietary models, ground-truth captions, or costly human annotations while maintaining robust temporal and spatial reasoning?
@@
To address this challenge, we propose VideoSAVi, a novel framework that requires no external supervision beyond the model itself. Our approach focuses on post-training refinement through a four-stage process: (1) generating diverse reasoning questions and initial answers about video content, (2) self-critiquing these answers to identify factual errors and reasoning flaws, (3) revising the responses based on the self-generated feedback, and (4) leveraging these pairs for DPO to align the model towards improved reasoning capabilities.
@@
Unlike prior approaches that depend on proprietary models or extensive annotations, VideoSAVi uniquely generates high-quality preference pairs through self-critique—identifying specific reasoning failures in both spatial relationships and temporal sequences. This self-supervision creates subtle yet unambiguous learning signals that, when optimized through DPO, enable efficient post-training refinement of capabilities already present in the model. By directly addressing the model’s own weaknesses rather than conforming to external judgment, our approach delivers significant improvements across diverse benchmarks without the computational cost of full retraining.
@@
    We introduce VideoSAVi, a novel self-training framework that enables Video-LLMs to reason over video content, eliminating the need for external supervision while maintaining computational efficiency.
@@
    We develop a quality-aware self-critique mechanism that generates preference pairs, focusing on both temporal and spatial reasoning, enabling comprehensive video understanding without external supervision.
@@
    Through extensive experiments, we demonstrate that VideoSAVi achieves state-of-the-art performance on MVBench (74.0%) and delivers significant improvements across multiple benchmarks, including +3.6% on NeXTQA, +3.9% on PerceptionTest, and +6.8% on EgoSchema. Moreover, VideoSAVi shows consistent effectiveness across various model architectures, with an average improvement of +3.4%.
@@ 2 Related Work
Recent Video-LLMs have made remarkable progress in multimodal understanding capabilities (Chen et al., 2024; Wang et al., 2024b; Li et al., 2024a; b; Zhang et al., 2024d). However, these models typically require massive instruction-tuning datasets and extensive computational resources for pre-training. Even with substantial training data, they often struggle with proper grounding in visual content (Fu et al., 2024; Zhou et al., 2024; Liu et al., 2024b), particularly for videos (Wang et al., 2024a; Yao et al., 2024; Wang et al., 2024e; Song et al., 2024). While some approaches attempt to address these challenges through architecture modifications (Liu et al., 2024a; Shu et al., 2024; Li et al., 2024d; Wang et al., 2024d), specialized training objectives (Lee et al., 2024; Lan et al., 2024; Liu et al., 2024c), or inference-time adaptations (Yang et al., 2024; Hu et al., 2025; Xu et al., 2024), they typically focus on specific aspects rather than holistic alignment. In contrast to these approaches that emphasize training from scratch or adding specialized components, VideoSAVi offers a post-training refinement strategy. Rather than attempting to teach new capabilities, we focus on identifying, amplifying, and improving specific reasoning skills already present in the model. This targeted enhancement allows us to efficiently boost performance in particular domains without the computational costs of full retraining. By leveraging self-alignment through preference optimization, VideoSAVi helps models rectify their own errors and strengthen existing capabilities across temporal and spatial dimensions without depending on external supervision or extensive new datasets.
@@
Recent work has adapted preference learning techniques to video-language models to improve alignment. LLaVA-Hound-DPO (Zhang et al., 2024b) introduced using Direct Preference Optimization (DPO) (Rafailov et al., 2023) for video understanding, but operates primarily at the text level without taking into account visual context. It requires preference pairs generated using proprietary models like GPT-4, introducing an external dependency. Similarly, Temporal Preference Optimization (TPO) (Li et al., 2025) focuses exclusively on temporal reasoning neglecting spatial relationships and requires video captioning as an intermediate step, adding computational overhead. Unlike these approaches that rely on preferences distilled from larger models or auxiliary tasks like captioning, VideoSAVi generates preference pairs directly from the model’s own assessment of right and wrong. This self-alignment approach targets both temporal and spatial dimensions simultaneously without external supervision. By having the model critique its own responses, we create a training signal that directly addresses the model’s specific failures rather than conforming to external judgement. This approach is more efficient and more focused on the model’s actual weaknesses rather than general quality improvements.
@@
Self-training has become a powerful method for improving language model performance (Gulcehre et al., 2023; Singh et al., 2024; Huang et al., 2023; Zelikman et al., 2022; Yeo et al., 2024; Wang et al., 2024c). The main idea is for models to generate their own training data and use it to refine their performance iteratively, with notable success in enhancing reasoning and task-specific capabilities. Recent advances have applied self-training to vision-language models (VLMs) (Sun et al., 2025; 2024; Deng et al., 2024) and Video-LLMs (Zohar et al., 2024). However, extending self-training to the video domain introduces unique challenges due to the complex spatiotemporal nature of videos. Existing approaches for Video-LLMs either rely on expensive teacher models (Sun et al., 2024), require ground truth labels for feedback (Zohar et al., 2024), or focus on single-aspect improvements. Furthermore, these methods often face limitations in generating high-quality preference data that effectively captures the relationships between visual elements across time (Yin et al., 2024). Our work addresses these challenges by introducing a novel self-alignment framework specifically designed for video understanding. We integrate self-critiquing mechanism with preference-based learning to enable Video-LLMs to automatically identify and improve upon their own reasoning errors in both spatial and temporal dimensions.
@@ 3 VideoSAVi
We present VideoSAVi (Figure 1), a novel self-aligned framework for enhancing video-language models without external supervision. Our approach leverages existing Video-LLMs to generate challenging questions, produce answers, critique responses, and refine them into high-quality outputs. This self-training pipeline addresses the high cost of human annotations and proprietary models and the need for comprehensive temporal and spatial reasoning.
Figure 1: Overview of VideoSAVi. Our iterative self-training pipeline consists of four key stages: (1) generating diverse spatial and temporal reasoning questions and initial responses about video content, (2) implementing a self-critique mechanism to identify reasoning errors and inconsistencies, (3) refining responses based on the critique feedback, and (4) training the model through DPO using paired initial (dispreferred) and revised (preferred) responses. This self-supervised cycle progressively improves VideoSAVi’s reasoning capabilities without external supervision.
VideoSAVi operates as a four-stage pipeline: (1) generating reasoning-focused questions and initial responses about video content, (2) implementing self-critique to identify spatial and temporal reasoning errors, (3) revising responses based on the critique feedback, and (4) optimizing model performance through DPO using the revised responses. The framework prompts the baseline model InternVL2.5 (Chen et al., 2024) along with the video to generate diverse questions about video content, focusing on temporal relationships and spatial details. The model then answers these questions, identifies reasoning errors through self-critique, and generates improved alternatives for optimization.
@@ 3.1 Question Generation and Initial Responses
For self-alignment to succeed, generated questions must target specific reasoning capabilities that the model can improve through preference optimization. We prompt the baseline model with an instruction template (detailed in Appendix, Figure 8) across two primary reasoning dimensions:
@@
We prompt the baseline model with instructions such as “Generate questions that require identifying spatial relationships between objects” and “Generate questions about visual details that might be easily overlooked.” The resulting questions focus on visual details, spatial relationships, and contrastive reasoning.
@@
We prompt the baseline model with instructions such as “Generate a question about the temporal sequence of events” and “Generate questions that require understanding what happened before or after specific key moments.” The resulting questions challenge the model to track objects across frames and understand cause-effect relationships.
Despite advancements in Video-LLMs, initial responses for these questions often contain reasoning errors in temporal sequencing, spatial relationships, and object detection. VideoSAVi addresses these challenges through a self-critiquing step where the model evaluates and refines its own reasoning without external supervision.
@@ 3.2 Self-Critique Mechanism
The self-critique process begins after the model produces an initial response $a_{0}$ to a question-video pair $(v,q)$. The model then acts as its own critic by prompting itself with a specialized template (detailed in Appendix, Figure 9) designed to elicit critical analysis:
@@
This critique function examines the answer for inconsistencies by comparing claims made in the response with visual evidence from the video frames. For spatial reasoning, the critique assesses object positions, visibility, and attributes, while for temporal reasoning, it analyzes event sequencing, causality, and transitions.
The critique prompt encourages the model to engage in counterfactual reasoning, considering what alternative interpretations of the video might be more accurate. This process is similar to the introspection mechanisms observed in human reasoning, where assessment of one’s own conclusions leads to refinement. The model learns to balance confidence in its initial assessment with appropriate skepticism, especially for ambiguous visual content.
@@
The critique identifies specific errors $\{e_{1},e_{2},...,e_{n}\}$ in the initial response. The severity assessment function $g_{\phi}$ is not explicitly learned through a separate training process, but rather is implicitly encoded in the language model’s parameters through the preference optimization process:
@@
This implicit severity assessment emerges from the model’s understanding of how different types of errors impact response quality, refined through the preference optimization process. Errors are categorized into critical (factual inaccuracies, logical contradictions) and minor (imprecisions, stylistic issues) based on their impact on overall correctness.
The critique deliberately targets common failure modes in video understanding, such as hallucination of non-existent objects or actions, temporal ordering errors that misrepresent the sequence of events, causal attribution errors that incorrectly identify relationships between actions, spatial relationship misinterpretations, and entity tracking failures across different frames.
@@ 3.3 Response Refinement
The refinement function is implemented easily through a prompt template rather than a separately parameterized function. The initial response $a_{0}$ and the generated critique $c$ are combined into a new prompt (detailed in Appendix, Figure 10) that instructs the model to produce an improved answer:
@@
This refinement prompt includes instructions like “Consider the following critique of your previous answer and provide an improved response that addresses these issues.” The model then generates a refined answer $a_{1}$ that incorporates the feedback from the self-critique.

### v3

@@ Abstract
Recent advances in video-large language models (Video-LLMs) have led to significant progress in video understanding. Current preference optimization methods often rely on proprietary APIs or human-annotated captions to generate preference data (i.e., pairs of model outputs ranked by quality or alignment with human judgment), which is then used to train models for video-language alignment. This approach is both costly and labor-intensive. To address this limitation, we introduce VideoSAVi (Self-Aligned Video Language Model), a self-training pipeline that enables Video-LLMs to learn from video content without external supervision. Our approach includes a self-critiquing mechanism that identifies reasoning errors in the model’s initial responses and generates improved alternatives, creating preference pairs directly from video content. VideoSAVi then applies Direct Preference Optimization (DPO) to iteratively train the model using the preference data, thus enhancing its temporal and spatial reasoning for video understanding. Experiments show that VideoSAVi delivers significant improvements across multiple benchmarks, including a +4.2 percentage point gain on MVBench, +3.9 on PerceptionTest, and +6.8 on the challenging EgoSchema dataset compared to baseline models.  Our model-agnostic approach is computationally efficient, requiring only 32 frames, offering a promising direction for self-aligned video understanding without reliance on external models or annotations.
@@ 1 Introduction
Vision-language models (VLMs) (Liu et al., 2023; Li et al., 2023a; Radford et al., 2021; Wang et al., 2024b; Chen et al., 2024; Li et al., 2024a) have made significant strides by integrating visual perception with the reasoning capabilities of large language models (LLMs) (OpenAI, 2024; Dubey et al., 2024; Ouyang et al., 2022). These models excel in interpreting and generating contextually relevant responses through the combination of image encoders and language generation techniques. Building on this foundation, recent video-large language models (Video-LLMs) (Zhang et al., 2023; Lin et al., 2024; Li et al., 2024b; Zhang et al., 2024d; Wang et al., 2022; 2024e) incorporate temporal information by converting video frames into tokens that LLMs can process, allowing them to better analyze video content. While Video-LLMs show impressive capabilities, they generally rely on large, high-quality annotated datasets, which are resource-intensive to obtain and limit scalability.
@@
Instruction tuning has been pivotal in advancing both VLMs and Video-LLMs (Liu et al., 2023; Brown et al., 2020; Xu et al., 2023; Wei et al., 2021; Wang et al., 2024b; Chen et al., 2024; Zhang et al., 2024c), but creating large-scale video instruction datasets is costly  (Deng et al., 2024). Recent efforts to generate massive instruction datasets (up to 1.3M pairs) by distilling knowledge from proprietary models like GPT-4V (Zhang et al., 2024d) have yielded only marginal improvements. This reliance on extensive annotated data and proprietary models limits the adaptability of Video-LLMs, creating barriers to wider adoption.
@@
Alignment-based approaches have shown promise for improving video understanding (Kulkarni & Fazli, 2025a). LLaVA-Hound (Zhang et al., 2024b) uses Direct Preference Optimization (DPO) (Rafailov et al., 2023) with text-based rankings from proprietary models, while TPO (Li et al., 2025b) focuses on temporal preference optimization by sampling negatives from video segments outside the target clip. However, both methods rely heavily on proprietary APIs or ground-truth captions, which limits their accessibility and scalability.
@@
This raises a critical research question: How can we train Video-LLMs to generate high-quality outputs without relying on proprietary models, ground-truth captions, or expensive human annotations while maintaining robust temporal and spatial reasoning?
@@
To address this challenge, we propose VideoSAVi, a novel framework requiring no external supervision beyond the model itself. Our approach focuses on post-training refinement through a four-stage process: (1) generating diverse reasoning questions and initial answers about video content, (2) self-critiquing these answers to identify factual errors and reasoning flaws, (3) revising the responses based on the self-generated feedback, and (4) leveraging these pairs for DPO to align the model towards improved reasoning capabilities.
@@
Unlike prior approaches that depend on proprietary models or extensive annotations, VideoSAVi uniquely generates high-quality preference pairs through self-critique, identifying specific reasoning failures in both spatial relationships and temporal sequences. This self-supervision creates subtle yet unambiguous learning signals that, when optimized through DPO, enable efficient post-training refinement of capabilities already present in the model. By directly addressing the model’s own weaknesses rather than conforming to external judgment, our approach delivers significant improvements across diverse benchmarks without the high computational cost of full retraining.
@@
    We introduce VideoSAVi, a novel self-training framework that enables Video-LLMs to reason over video content without external supervision, while maintaining computational efficiency.
@@
    We develop a quality-aware self-critique mechanism that generates preference pairs by focusing on both temporal and spatial reasoning, enabling comprehensive video understanding.
@@
    Through extensive experiments, we demonstrate that VideoSAVi delivers significant improvements across multiple benchmarks, including +4.2 percentage points on MVBench, +3.6 on NeXTQA, +3.9 on PerceptionTest, and +6.8 on EgoSchema. Moreover, VideoSAVi consistently improves various model architectures, with an average gain of +2.1 percentage points across all benchmarks.
@@ 2 Related Work
Recent Video-LLMs have made remarkable progress in multimodal understanding (Chen et al., 2024; Wang et al., 2024b; Li et al., 2024a; b; Zhang et al., 2024d). However, these models typically require massive instruction-tuning datasets and extensive computational resources for pre-training. Even with substantial training data, they often struggle with proper grounding in visual content (Fu et al., 2024; Zhou et al., 2024; Liu et al., 2024b; Kulkarni & Fazli, 2025b), particularly for videos (Wang et al., 2024a; Yao et al., 2024; Wang et al., 2024e; Song et al., 2024). While some approaches attempt to address these challenges through architecture modifications (Liu et al., 2024a; Shu et al., 2024; Li et al., 2024d; Wang et al., 2024d), specialized training objectives (Lee et al., 2024; Lan et al., 2024; Liu et al., 2024c), or inference-time adaptations (Yang et al., 2024; Hu et al., 2025; Xu et al., 2024), they typically focus on specific aspects rather than holistic alignment. In contrast, VideoSAVi introduces a post-training refinement strategy that avoids training from scratch or adding complex components. Instead of introducing entirely new capabilities, our method identifies, amplifies, and improves reasoning skills already present in the model. By leveraging self-alignment through preference optimization, VideoSAVi enables models to correct their own errors and strengthen existing temporal and spatial reasoning, without relying on external supervision or large new datasets.
@@
Recent work adapts preference learning techniques to video-language models to improve alignment. LLaVA-Hound-DPO (Zhang et al., 2024b) applies Direct Preference Optimization (DPO) (Rafailov et al., 2023) to video understanding, but operates mainly at the text level, without incorporating visual context. It relies on preference pairs from proprietary models like GPT-4, introducing external dependencies. Similarly, Temporal Preference Optimization (TPO) (Li et al., 2025b) focuses solely on temporal reasoning, ignores spatial relationships, and requires video captioning as an intermediate step, adding computational overhead. In contrast, VideoSAVi generates preference pairs directly from the model’s own assessment of correctness, eliminating the need for external supervision or auxiliary tasks like captioning. This self-alignment approach simultaneously targets both temporal and spatial reasoning. By critiquing its own responses, the model produces training signals that directly address its specific reasoning failures rather than conforming to external judgments. This results in a more efficient and focused refinement process, targeting the model’s actual weaknesses rather than pursuing general improvements.
@@
Self-training is a powerful method for improving language model performance (Gulcehre et al., 2023; Singh et al., 2024; Huang et al., 2023; Zelikman et al., 2022; Yeo et al., 2024; Wang et al., 2024c). The core idea is for models to generate their own training data and use it to refine their performance iteratively, with notable success in enhancing reasoning and task-specific capabilities. Recent advances extend self-training to vision-language models (VLMs) (Sun et al., 2025; 2024; Deng et al., 2024) and Video-LLMs (Zohar et al., 2024). However, applying self-training to the video domain introduces unique challenges due to the complex spatiotemporal nature of videos. Existing methods often rely on expensive teacher models  (Sun et al., 2024), require ground truth labels (Zohar et al., 2024), or focus on isolated aspects of video understanding. Moreover, they struggle to generate high-quality preference data that captures nuanced relationships between visual elements across time (Yin et al., 2024). Our work addresses these limitations by introducing a novel self-alignment framework tailored for video understanding. We integrate a self-critiquing mechanism with preference-based learning, enabling Video-LLMs to automatically detect and correct their own reasoning errors across spatial and temporal dimensions.
@@ 3 VideoSAVi
We present VideoSAVi, a novel self-aligned framework that enhances video-language models without relying on external supervision. VideoSAVi leverages existing Video-LLMs to generate challenging questions, produce and critique answers, and refine them into high-quality outputs. This self-training pipeline addresses the high costs associated with human annotations and proprietary models, while also improving the model’s temporal and spatial reasoning capabilities. VideoSAVi follows a four-stage process: (1) generating reasoning questions and initial responses based on video content, (2) identifying temporal and spatial reasoning errors through self-critique, (3) revising responses based on the critique, and (4) optimizing model performance using DPO with the improved responses. Specifically, the framework prompts the baseline model, InternVL2.5 (Chen et al., 2024), along with the video input, to produce diverse questions that probe complex temporal dynamics and spatial details. The model then answers these questions, critiques its own reasoning, and generates refined responses to guide further learning. Figure 1 illustrates the overall architecture and iterative self-training pipeline of VideoSAVi.
Figure 1: Overview of VideoSAVi. Our iterative self-training pipeline consists of four key stages: (1) generating diverse spatial and temporal reasoning questions and initial responses about video content, (2) applying a self-critique mechanism to identify reasoning errors and inconsistencies, (3) refining responses based on the critique feedback, and (4) training the model through DPO using paired initial (dispreferred) and revised (preferred) responses. This self-supervised cycle progressively improves VideoSAVi’s reasoning capabilities without external supervision.
@@ 3.1 Question Generation and Initial Responses
For self-alignment to be effective, generated questions must target specific reasoning skills that can be refined through preference optimization. To this end, we prompt the baseline model using an instruction template (see Appendix, Figure 8) across two core reasoning dimensions:
@@
We use prompts such as “Generate questions that require identifying spatial relationships between objects” and “Generate questions about visual details that might be easily overlooked.” The resulting questions focus on spatial layouts, fine-grained visual details, and contrastive reasoning.
@@
We use prompts such as “Generate a question about the temporal sequence of events” and “Generate questions that require understanding what happened before or after specific key moments.” The resulting questions challenge the model to track objects across frames and infer cause-and-effect relationships.
Despite advancements in Video-LLMs, their initial responses to such questions often reveal reasoning errors in spatial relationships, temporal sequencing, and object tracking. VideoSAVi addresses these limitations through a self-critiquing mechanism that allows the model to evaluate and improve its own reasoning without external supervision.
@@ 3.2 Self-Critique Mechanism
The self-critique process begins once the model generates an initial response $a_{0}$ to a question-video pair $(v,q)$. It then assumes the role of its own critic by prompting itself with a specialized template (see Appendix, Figure 9) designed to elicit a critical evaluation of its reasoning:
@@
The critique function analyzes the initial answer for inconsistencies by comparing its claims against visual evidence from the video frames. For spatial reasoning, it evaluates object positions, visibility, and attributes. For temporal reasoning, it examines event sequencing, causality, and transitions. The critique prompt encourages counterfactual reasoning by asking the model to consider alternative interpretations that may better align with the video content. This self-assessment mirrors human introspection, where evaluating one’s own conclusions leads to refinement. Through this process, the model learns to balance confidence in its initial assessment with appropriate skepticism, particularly in the face of ambiguous visual input.
@@
The critique identifies specific errors $\{e_{1},e_{2},...,e_{n}\}$ in the initial response. Rather than being learned through a separate training process, the severity assessment function $g_{\phi}$ is implicitly encoded in the model’s parameters via preference optimization:
@@
The model implicitly learns to assess error severity by developing an understanding, refined through preference optimization, of how different types of mistakes affect response quality. It categorizes errors as either critical (e.g., factual inaccuracies, logical contradictions) or minor (e.g., imprecisions, stylistic issues) based on their influence on overall correctness.
The critique specifically targets common failure modes in video understanding, including hallucinations of non-existent objects or actions, errors in temporal ordering that misrepresent event sequences (Li et al., 2025a), incorrect causal attributions, misinterpretations of spatial relationships, and failures in tracking entities across frames.
@@ 3.3 Response Refinement
The refinement function is implemented via a prompt template rather than a separately parameterized module. The initial response $a_{0}$ and the generated critique $c$ are combined into a new prompt (see Appendix, Figure 10) that instructs the model to produce an improved answer:
@@
This refinement prompt includes instructions such as “Consider the following critique of your previous answer and provide an improved response that addresses these issues.” The model then generates a refined answer $a_{1}$ that incorporates the feedback from the self-critique.